Textual Variation of the New Testament

Matthew Gray

March 13, 1997

Introduction

Many documents, such as the works of Shakespeare, have had histories of questions as to their textual consistency from one version to another. Questions have been raised as to which version is the “original” or “true” version, whether the accepted author is the true author, and whether these questions are meaningful on their own.

While this sort of debate is common regarding literature and historical documents, it is somewhat more taboo when discussing religious documents, such as the New Testament of the Bible. It becomes a particularly difficult issue in the case of religious documents which claim to be the infallible word of a god.

The question remains: How much variation exists in the texts of the New Testament?

Types of Variations

There are three primary types of variations, two of which will be discussed here. These types are internal variation, manuscript variation, and translation variation.

Internal variation, or so-called “Bible Contradictions” are the type that are not to be discussed here. While it is apparent that these exist plentifully from the obvious such as Jesus’ lineage in Matthew and Luke, to the substantially more debatable and subtle, they are in a seperate class and should be discussed seperately.

Manuscript variation, or variation from one copy of the New Testament to another is substantially more akin to the kinds of issues often associated with literary or historical documents. This type of variation will be the focus of this paper.

Translation variation, or variations between translations of the same source document, is also relevent. Additionally, issues of manuscript variation come into this, as individual translations may use different source manuscripts.

Manuscript variation

There are no manuscripts of the New Testament, allegedly a first century document, dating from the first century. Thousands of Greek manuscripts exist, dating from the 7th century, and a few dozen from before that. Substantial variations exist between these manuscripts, ranging from obvious copyist errors to genuine changes in content. Analysis and organization of these variations abounds.^{1} More so, what documents are included in the “New Testament” varies quite a bit among these, often excluding some books, or sections of books, or including books that are no longer considered part of the New Testament. One theory is that all of these manuscripts originated from some first century document dubbed “Documenta Q”,^{2} though no versions of this document have been found.

1http://www.skypoint.com/ waltzmn/Examples.html

Four manuscripts from the fourth century are among the most widely used sources. Only one of these, Codex Sinaiticus, contains the entire New Testament, and it includes additional material other than the New Testament. The other three manuscripts are: Codex Vaticanus, which is missing the end of the New Testament, Codex Alexandrinus, which is missing much of Matthew and some of John, and Codex Bezae, which contains Matthew, John, Luke, Mark, and Acts.^{3}

Though there is extensive specific variation in content of these manuscripts the general content remains impressively consistent for a roughly 2 millenia old document. Which specific books were included in the New Testament, however, seems to be a more careful selection of medieval and later scholars, rather than any first or even second century intent.

Translation variation

A great many translations of the New Testament exist, in English and other languages. Popular English versions include the King James Version, the New International Version, the American Standard Version, the Revised Standard Version and others.^{4}.

Among these translations, different interpretations have been made of many parts, and it is difficult, if not impossible, to know the original meaning. Though old manuscripts do exist, languages change over centuries and the sense of words, or use of idiom drastically affects a translation.

Technological Impact

With the advent of the printing press, the consistency of the New Testament increased dramatically. No longer were manuscripts subject to the errors or editorial whims of the copyists, as they had been previously. This substantially decreased, if not altogether stopped manuscript variations, but it left the issue of which manuscript to use as source.

Presently, the near total ubiquity of printed versions of the New Testament creates a perception of a textually invariant Bible, with the possible exception of translation variation. Popularly, however, translation variations are perceived to be primarily in style of language (“was the father of” v. “begat”), rather than actual content. This, combined with social taboos against questioning textual reliability of the Bible, have created the illusion of a perfect, unchanging text.

On the other hand, electronic versions of the Bible^{5} have made contrasting translations substantially easier. Similar work has been done for comparison of manuscripts.^{6}

Conclusion

The New Testament, like many literary and historical documents, has undergone substantial textual variation. Though, like the plays of Shakespeare, there may be substantial variation in specifics, for the most part, the broad message is fairly consistent across manuscripts and translations.

2http://www.theatlantic.com/atlantic/issues/96dec/jesus/jesus.htm

3http://www.bowness.demon.co.uk/reli0001.htm

4http://www.gospelcom.net/bg/translations.html

5http://www.gospelcom.net/bible

6http://www.entmp.org/entmp/papers/coll.html

Recent technologies, including the printing press, have both created a facade of constancy, as well as enabled analysis of the variations that do exist.